OpenStudy (anonymous):

.If ON = 9x – 4, LM = 8x + 7, NM = x – 3, and OL = 4y – 8, find the values of x and y for which LMNO must be a parallelogram. A. x = 11, y = 4 B. x = 11, y = 11 C. x = 4, y = 4 D. x = 4, y = 11

OpenStudy (anonymous):

Well, as we know, a parallelogram has 2 equal sides of one length and 2 equal sides of another. As such, we find the lines which to not touch, such as ON and LM for one group and NM and OL for another, and make them equal, solving for the required variable. ON = LM 9x - 4 = 8x + 7 x - 4 = 7 x = 11 NM = OL x - 3 = 4y - 8 Since we already found x, we just plug it in and: 11 - 3 = 4y - 8 8 = 4y - 8 16 = 4y 4 = y

OpenStudy (anonymous):

other word its a